What is the unit rate of 84 pages read in 16 minutes
Answer by tommyt3rd(5050)

16 min =16/60 hr = 4/15 hr
84 pages in 4/15 hr is

(84/1)/(4/15)=
84*(15/4)=
315 pages per minute